How many 2 by 2 by 2 cubes must be added to an 8 by 8 by 8 cube to make a 10 by 10 by 10 cube ?

Answer by Alan3354(69443)

10^3 = 1000
8^3 = 512
------------ Subtract
488
2^3 = 8, divide by 8
---------
488/8 = 61 cubes